OpenStudy (anonymous):

@jojo4eva What is the first step to solve the equation b minus five equals eight? Add 8 to both sides. Subtract 8 from both sides. Add 5 to both sides. Subtract 5 from both sides.

OpenStudy (mathmale):

Rose: What do YOU think?

OpenStudy (jojo4eva):

Add 5 to both sides rose

OpenStudy (jojo4eva):

when u add 5 to both sides tou then get b=13

OpenStudy (anonymous):

b - 5 = 8 + 5 +5 b = 13 So you can tell us what the first step is by looking at how I solved this?

OpenStudy (anonymous):

adding five at both sides @Victoria_Davis_2018

OpenStudy (anonymous):

correct....add 5 to both sides is the first step

OpenStudy (anonymous):

Yes Rose

OpenStudy (mathmale):

Rose, what about you? What do YOU think?